OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of each should he use? Let p = pounds of peanuts and let c = pounds of cashews. Write a system of equations that could be used to solve the problem.

OpenStudy (anonymous):

@Elijah16

OpenStudy (anonymous):

OK i know one equation would be 4x+6.5y=4.75 but what is the second equation

OpenStudy (anonymous):

@a1234

OpenStudy (a1234):

The other equation is p + c = 10.

OpenStudy (anonymous):

P = 7 C = 3

OpenStudy (anonymous):

thanks both of you

OpenStudy (a1234):

You're welcome.

OpenStudy (anonymous):

you are super helpful

OpenStudy (anonymous):

Write an equation for the total weight. p + c = 10 Write an equation for the mixture. Remember the final mixture is $4.75 for each of 10 lbs so multiply 4.75 x 10: 4.00p + 6.50c = 4.75(10) Let's simplify the second equation: 4p + 6.5c = 47.5 Multiply both sides by 2 to get rid of the decimals: 8p + 13c = 95 Then solve the first equation for one variable (say p): p = 10 - c Substitute: 8(10-c) + 13c = 95 Distribute the 8 through the parentheses: 80 - 8c + 13c = 95 80 + 5c = 95 Subtract 80 from both sides: 5c = 15 Divide both sides by 5: c = 15/5 c = 3 Then substitute back to get p: p = 10 - c p = 10 - 3 p = 7 Answer: 7 lbs of peanuts 3 lbs of cashews Source: Yahoo